# 2]organization=Shanghai Blockchain Engineering Research Center, city=Shanghai, postcode=200433, country=China

### v1

@@ 6.4.1 Size of the auxiliary dataset
We tested four sizes of auxiliary datasets: 0.0025, 0.025, 0.125, and 0.25, representing their relative sizes to the VFL training dataset. Results in Figure 5 show that as the auxiliary dataset size increases, the reconstruction accuracy of our method improves. This suggests that a larger auxiliary dataset, offering more information, allows for a more accurate estimation of the original dataset’s distribution, thus enhancing reconstruction accuracy.
@@ 6.4.3 Size of the Known Private Dataset
In the SA attack scenario, we assume that the attacker has acquired a small number of the target’s private data to train an InverNet for data reconstruction. We evaluated the attack’s efficacy with varying numbers of prior-known private samples, 8, 16, 32, and 64, as shown in Figure 6 and Table 6.
@@ 6.4.4 The Impact of InverNet Model Size on Attack Effectiveness
6.5 Defense evaluation
Bank Adult
Ratio AUC QA DPA IQA SA Ratio AUC QA DPA IQA SA
$1$ $85.89$ $86.38$ $88.29$ $18.96$ $18.96$ $1$ $88.79$ $87.78$ $84.92$ $73.92$ $73.87$
$0.5$ $93.80$ $97.46$ $95.40$ $52.80$ $18.96$ $0.5$ $87.71$ $92.27$ $87.59$ $84.15$ $83.07$
$0.1$ $93.84$ $97.93$ $95.50$ $64.01$ $78.16$ $0.1$ $87.62$ $93.51$ $89.29$ $82.91$ $81.45$
$0.01$ $93.81$ $98.31$ $96.02$ $69.68$ $89.48$ $0.01$ $86.95$ $95.44$ $88.56$ $82.94$ $87.50$
$0.001$ $93.92$ $98.22$ $96.04$ $66.40$ $90.49$ $0.001$ $86.66$ $95.86$ $89.15$ $61.72$ $85.48$
Credit CIFAR10
Ratio AUC QA DPA IQA SA Ratio ACC QA DPA IQA SA
$1$ $77.03$ $96.75$ $92.87$ $28.92$ $88.46$ $1$ $61.99$ $0.70$ $0.85$ $0.34$ $0.11$
$0.5$ $76.02$ $96.52$ $96.52$ $26.04$ $90.45$ $0.5$ $67.26$ $0.14$ $0.81$ $0.14$ $0.0$
$0.1$ $77.36$ $97.95$ $96.00$ $39.92$ $93.13$ $0.1$ $72.99$ $0.72$ $0.87$ $0.44$ $0.68$
$0.01$ $77.16$ $97.59$ $96.25$ $41.50$ $93.26$ $0.01$ $74.84$ $0.80$ $0.88$ $0.53$ $0.74$
$0.001$ $76.48$ $98.36$ $95.09$ $41.05$ $93.29$ $0.001$ $74.25$ $0.86$ $0.89$ $0.57$ $0.81$
Bank Adult
Ratio AUC QA DPA IQA SA Ratio AUC QA DPA IQA SA
$1$ $93.68$ $18.96$ $89.19$ $49.02$ $18.96$ $1$ $88.53$ $70.87$ $69.93$ $56.21$ $3.07$
$0.5$ $93.94$ $98.52$ $94.51$ $64.81$ $18.96$ $0.5$ $89.18$ $91.20$ $71.88$ $90.92$ $3.07$
$0.1$ $94.01$ $98.22$ $95.48$ $67.16$ $90.67$ $0.1$ $87.02$ $96.37$ $90.68$ $73.40$ $87.19$
$0.01$ $93.95$ $97.85$ $95.59$ $63.60$ $88.81$ $0.01$ $88.55$ $91.14$ $82.20$ $72.69$ $83.50$
$0.001$ $93.85$ $98.26$ $96.20$ $60.42$ $91.19$ $0.001$ $86.81$ $94.90$ $89.17$ $78.52$ $86.16$
Credit CIFAR10
Ratio AUC QA DPA IQA SA Ratio ACC QA DPA IQA SA
$1$ $76.90$ $98.13$ $87.77$ $62.98$ $3.46$ $1$ $64.97$ $0.75$ $0.79$ $0.59$ $0.0$
$0.5$ $77.51$ $98.35$ $96.08$ $57.51$ $94.34$ $0.5$ $66.81$ $0.88$ $0.79$ $0.62$ $0.07$
$0.1$ $76.92$ $98.04$ $96.20$ $33.98$ $91.86$ $0.1$ $74.41$ $0.86$ $0.83$ $0.59$ $0.82$
$0.01$ $77.23$ $97.89$ $96.24$ $37.66$ $91.96$ $0.01$ $74.13$ $0.81$ $0.87$ $0.56$ $0.76$
$0.001$ $76.81$ $97.54$ $95.35$ $33.91$ $92.52$ $0.001$ $74.38$ $0.80$ $0.90$ $0.57$ $0.90$
Table 9: Experimental Results of Gaussian noise Defense. For each dataset, the first column represents the ratio of the defense, the second column shows the results of the VFL task, and the last four columns indicate the effectiveness of our attack method under four different scenarios. For CIFAR10, SSIM is the Evaluation Metric for the Last Four Columns.
Although this study primarily focuses on attack strategies, we have also examined several defensive measures. Particularly, effective defense methods such as Differential Privacy (DP), noise addition, purification defense strategies Yang et al. (2023), and the adaptive obfuscation privacy protection strategy FedPass Gu et al. (2023) have been considered. Differential privacy technology demonstrates significant potential in defending against data reconstruction attacks by adding noise to gradients during the training process, thereby protecting individual data privacy while maintaining the overall effectiveness of the model. The method of adding noise to model outputs hinders attackers from obtaining precise information, thus protecting the data from malicious use, although this may impact the accuracy of the model. Purification defense strategies use technologies like Autoencoders to reduce the variability in model outputs, enhance privacy protection, and decrease the distinguishability between outputs, thereby reducing the likelihood of attackers using model outputs for inference. FedPass, as a novel defense framework, is particularly suitable for Vertical Federated Learning (VFL) environments, employing adaptive obfuscation techniques to protect data privacy while allowing contributions from various parties.
To assess the robustness of our UIFV method, we implemented two representative data reconstruction defense methods within the VFL framework, namely Differential Privacy and Gaussian noise addition, to test the performance of UIFV. In our experiments, we set the noise ratio to 1, 0.5, 0.1, 0.01, and 0.001, respectively, and conducted experiments across four different scenarios in four datasets. The experimental results of Differential Privacy are shown in Table 8, and those of noise addition are in Table 9.
Overall, these two defense methods do indeed have an effect in reducing the efficacy of data reconstruction attacks. However, the impact of these methods on the effectiveness of attacks is relatively limited. In scenarios such as QA, DPA, and IQA, a certain success rate of attacks can still be maintained even with a high noise ratio. In the SA scenario, the impact of these two defense methods is more noticeable. As the noise ratio increases, there is a downward trend in the attack accuracy in the SA scenario.
7 Conclusion
@@
In our paper, we introduce the Unified InverNet Framework in VFL (UIFV) a novel approach for conducting data reconstruction attacks in VFL environments. This method diverges from traditional attack strategies by utilizing the intermediate features of the target model instead of relying on gradient information or model information. UIFV demonstrates exceptional adaptability, suitable for various black-box scenarios. Experiments conducted on four benchmark datasets show that our approach surpasses the existing attack methods in effectiveness, achieving over $96\%$ accuracy in scenarios like QA. Through in-depth ablation studies, we also confirm the significance of key components such as the data generation module. This research highlights the significant privacy risks faced by VFL systems and underscores the urgent need to establish robust defense mechanisms to counter such attacks.
8 CRediT authorship contribution statement
Jirui Yang: Methodology, Writing - original draft. Peng Chen: Methodology, Writing - review and editing. Zhihui Lu: Methodology, Supervision, Writing- review and editing. Qiang Duan: Methodology, Writing - review and editing. Yubing Bao: Formal analysis, Data curation.
9 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 Acknowledgements
The work of this paper is supported by the National Key Research and Development Program of China (2022YFC3302300, 2021YFC3300600), National Natural Science Foundation of China under Grant (No. 92046024, 92146002, 61873309), and Shanghai Science and Technology Innovation Action Plan Project under Grant (No.22510761000).
figure[<options>] [Uncaptioned image]
Table 10:
[…]

### v2

@@ 6.5.1 Size of the auxiliary dataset
We tested four sizes of auxiliary datasets: 0.0025, 0.025, 0.125, and 0.25, representing their relative sizes to the VFL training dataset. Results in Fig. 8 show that as the auxiliary dataset size increases, the reconstruction accuracy of our method improves. This suggests that a larger auxiliary dataset, offering more information, allows for a more accurate estimation of the original dataset’s distribution, thus enhancing reconstruction accuracy.
@@ 6.5.3 Size of the Known Private Dataset
In the SA attack scenario, we assume that the attacker has acquired a small number of the target’s private data to train an InverNet for data reconstruction. We evaluated the attack’s efficacy with varying numbers of prior-known private samples, 8, 16, 32, and 64, as shown in Fig. 9 and Table 13.
@@ 6.5.4 The Impact of InverNet Model Size on Attack Effectiveness
7 Conclusions
7.1 Summary
@@
In our paper, we introduced the Unified InverNet Framework in VFL (UIFV), a novel approach for conducting data reconstruction attacks in VFL environments. Unlike traditional attack strategies, UIFV leverages intermediate features of the target model rather than relying on gradient information or model parameters. UIFV exhibits remarkable adaptability and is effective across various black-box scenarios. Experiments conducted on four benchmark datasets show that our approach surpasses the existing attack methods in effectiveness, achieving over $96\%$ accuracy in scenarios like QA. Through comprehensive ablation studies, we also confirmed the importance of key components, such as the data generation module. Our research expands the understanding of VFL data reconstruction attacks and provides new insights for privacy protection. The UIFV framework showcases its high applicability and precision across multiple scenarios, offering practical guidance for designing more robust defense mechanisms in the future. Additionally, our findings highlight privacy vulnerabilities in VFL systems under real-world applications, providing valuable support for policy-making and technological advancements in data protection.
7.2 Limitation
Despite its strong experimental performance, the UIFV framework has certain limitations. First, this study primarily focuses on scenarios where the attacker is an active participant in the VFL system. However, when the attacker acts as a passive participant, the conditions for a successful attack become significantly more stringent. For instance, in Data Passive Attack (DPA) scenarios, passive participants may find it challenging to train effective shadow models, potentially limiting the applicability of the UIFV method. Furthermore, our research is based on the most generic VFL architectures. More complex VFL setups could introduce additional challenges for the UIFV framework.
7.3 Future work
We consider adapting the UIFV framework to accommodate more complex VFL architectures and extend its application to real-world scenarios in healthcare, finance, and the Internet of Things (IoT). These efforts will help validate the framework’s broad applicability and practical impact. Additionally, considering the privacy risks exposed by the UIFV framework, future research should focus on developing more advanced defense mechanisms, such as purification defense strategies [40], to comprehensively enhance the security of VFL systems.